OpenStudy (welshfella):

Integrate 1 / (x^10 + x) 1

OpenStudy (welshfella):

I'm not sure about the best way to approach this one.

OpenStudy (welshfella):

partial fractions does n't help

ganeshie8 (ganeshie8):

multiply top and bottom by \(x^{-10}\) : \[\int \dfrac{x^{-10}}{1+x^{-9}}\,dx\]

OpenStudy (welshfella):

ah!!!

ganeshie8 (ganeshie8):

next step is obvious sub \(u = 1+x^{-9}\)

OpenStudy (welshfella):

so its -1/9 ln (1 + x^-9)?

ganeshie8 (ganeshie8):

looks good !

OpenStudy (welshfella):

thank you

ganeshie8 (ganeshie8):

np

OpenStudy (welshfella):

integration is an art...

ganeshie8 (ganeshie8):

definitely almost every integration problem requires guessing and makes mathematicians feel humble of the fact that we can't integrate everything into an elementary function ;)

OpenStudy (welshfella):

too true